if $8000 is invested in a bank that pays 4% interest, how much more must be invested in bonds at 9% to earn 5% on the total investment?
Answer by richwmiller(17219)

x+y=8000
.04x+.09y=.05(8000)
.04x+.09y=400
x=8000-y
0.04*(8000-y)+0.09*y=400
320-0.04y+0.09*y=400
0.05*y=80
y=1600 at 9%
x=8000-y
x=6400 at 4%
check
0.04*6400+0.09*1600=400
256+144=400
400=400
ok